# Re-Introducing LayerNorm: Geometric Meaning, Irreversibility and a Comparative Study with RMSNorm

### v1

@@ 4 Experiments
After a detailed theoretical analysis of the effects of layer normalization on a vector, we next study how LayerNorm affects the hidden representations of LLMs in practice. While prior works have studied the effects of LayerNorm on downstream model performance and training convergence (Ba et al., 2016; Xu et al., 2019; Zhang & Sennrich, 2019), there has been a surprising gap in the literature about studying the effect of LayerNorm on the internal representations of a model at inference time. To do so, we pass one million tokens from Wikipedia articles through various models and study how LayerNorm modifies a hidden vector in the representation space.
Methods and Models
We study the impact of LayerNorm on the hidden representations of 7 decoder-only LLMs across two size categories. The models used are listed in Table 1. GPT2XL, GPT-Neo-1.3B, and Pythia-1.4B represent the small LLM size category, whereas GPT-J-6B, Pythia-6.9B, Llama-2-7B, and Llama-3-8B represent the medium LLM size category. We pass one million tokens from Wikipedia articles through each model and capture the hidden representations for all tokens before and after each normalization layer. This is done for each layer inside the model, which creates many terabytes to be analyzed. Norms were calculated using the L2 norm, and angle differences in degrees were determined through cosine similarity calculations to measure the orientation changes induced by normalization.
As outlined in Table 1, we also use two types of normalization methods. Models like GPT-2 XL, GPT-Neo 1.3B, Pythia-1.4B, GPT-J 6B, and Pythia-6.9B employ LayerNorm (Ba et al., 2016), whereas Llama-2-7B and Llama-3-8B utilize RMSNorm (Zhang & Sennrich, 2019). The crucial difference between LayerNorm and RMSNorm is the subtraction of the mean vector (Figure 2(a)). In this paper, we study how LayerNorm and RMSNorm operate on the hidden representations.
@@
Figure 3: This figure shows the growing norm of the residual stream or hidden vectors at each layer (a-c) and how LayerNorm and RMSNorm regulate the growing norms (d-f) for GPTJ 6B, Pythia 6.9B and Llama3 8B.
@@ 4.1 Norm Stabilization
As a token representation passes through the different layers inside an LLM, the hidden representations accumulate due to residual connections (He et al., 2016). These hidden representations get added to the residual stream at each layer as shown in equation 6 and can cause the norm of the hidden vectors to grow. In fact, the norm of the hidden vectors at each layer grows disproportionately with large standard deviations, as can be seen in Figure 3 (a-c). The growing norm of the hidden vectors is extreme for some models more than others, as is seen for GPT-J and Pythia-6.9B. The growing norm of hidden representations for the remaining models from Table 1 and the effect of layer normalization can be seen in Figure 6 (appendix).
The significant role of LayerNorm and RMSNorm in stabilizing the growing norm of the hidden vectors at each layer can be seen in Figure 3 (d-f), which shows the norm of the hidden vectors after the first normalization step. This shows that normalization before sending vectors into the attention module significantly attentuates the vectors norms and reduces their variance. The standardization step in layer normalization first modifies the norm of each vector to $\sqrt{d}$, where $d$ is the dimensionality of the representation space. The norms of the standardized vectors are further modified marginally by the scale-and-shift steps. Layer normalization effects both the mean and the spread of the norms, where the standard deviations in norm reduce by factors between 10-100 depending on the model (Table 3 in appendix). These results show the crucial role that LayerNorm and RMSNorm play in stabilizing the intermediate hidden vectors in practice.
@@ 4.2 Rotation
While stabilizing the growing norm of hidden vectors is a major function of layer normalization, it also ends up rotating the hidden vectors. Figure 4 shows the angle between the original hidden vector and the post-normalization vector. As can be seen in Figure 4 (and in Figure 8 in appendix for the rest of the models), each layer is responsible for a fixed, non-trivial amount of rotation of hidden vectors in the representation space. Mean rotation angles in degrees between hidden and post-layer normalization vectors (LN1 and LN2) across all layers for various models are shown in Table 4 in the appendix. Thus, both LayerNorm and RMSNorm additionally rotate hidden vectors in representation space in addition to stabilizing the norms of the vectors.
4.3 LayerNorm versus RMSNorm : The need for “Mean Subtraction”
Apart from norm stabilization and rotation, a critical aspect of LayerNorm is its ability to orient the hidden vectors orthogonal to the uniform vector, as discussed in section 3. This is done by permanently (section 3.3) removing the component of every vector along the uniform vector and then normalizing the remaining vector. Such an action trains the model to operate orthogonal to the uniform vector, implying that the information along the uniform vector is not useful (section 3.4). RMSNorm, on the other hand, does no such thing and simply normalizes the existing vector, showing similar performance (Zhang & Sennrich, 2019).  Thus, a natural question to ask here is: “Is the mean subtraction step really needed?”
(a) GPT-J
(b) Pythia 6.9
(c) Llama-3
@@ 4.3 LayerNorm versus RMSNorm : The need for “Mean Subtraction”
Figure 5: Distribution of angles (in degrees) between Hidden vectors (a-c) and post-normalization vectors (d-f) with the uniform vector for GPT-J, Pythia 6.9, Llama-3 for a layer picked at random (Layer 24).
The Llama model series (Touvron et al., 2023a; b) is the most prominent family of LLMs that use RMSNorm. While the fact that Llama models achieve state-of-the-art results across multiple measures (Dubey et al., 2024) shows that using RMSNorm instead of LayerNorm does not hurt performance, it is hard to isolate the role of the “mean subtraction” step in overall performance. In this paper, we answer this question mechanistically by finding justification for the need for such an operation. To do so, we find the angle between the hidden vector and the uniform vector just before and after layer normalization operations (equations 1 and 3). The results are shown in Figure 5 for GPT-J (6B), Pythia (6.9B), and Llama-3 (8B). We see that for all models in Figure 5, the angle between the hidden vectors and the uniform vectors is 90 degrees on average even before normalization, with a very small spread. This remains true post-normalization as well for all three models.
@@
These results are surprising for two reasons. Firstly, the hidden representations for LayerNorm-based models (GPT-J and Pythia) are themselves orthogonal to the uniform vector even before the LayerNorm operation. This is true for 4 out of the 5 LayerNorm-based models studied in this paper except GPT2-XL (Figure 10 in the appendix). This shows that the “mean subtracting” operation of LayerNorm is redundant during inference. While indicating the redundancy of removing the component along the uniform vector during inference, it may be that the model learns to align the hidden representations orthogonal to the uniform vector during the training process, hence requiring the mean subtraction step. This is what makes the second observation even more surprising. We find that even for Llama-3 (8B) trained using RMSNorm, the hidden representations before normalization are also on average orthogonal to the uniform vector with a very small spread. In fact, they are more orthogonal to the uniform vector than for LayerNorm-based models, with a much smaller variance. The large variance for LayerNorm-based models is due to a long tail omitted in the figure. After RMSNorm, the hidden representations continue to be orthogonal to the uniform vector. This is also true for Llama-2 (7B) (Figure 10 in the appendix).
@@
These results show that the “mean subtracting” operation in LayerNorm is redundant as the model learns to operate orthogonal to the uniform vector naturally during the training process, as seen in Llama-3 (8B) and Llama-2 (7B) trained with RMSNorm and for 4 out of 5 models trained using LayerNorm. The model naturally aligns its representations orthogonal to the uniform vector as expected in such high-dimensional spaces and, as a consequence, promotes diversity in hidden vector components. With this, we provide the first mechanistic evidence that the “mean subtraction” step in LayerNorm is dispensable, as we do not find hidden vectors having large components along the uniform vector. Moreover, the irreversibility of LayerNorm means that the information lost in the “mean subtraction” step is never recovered. Due to this, we advocate for using RMSNorm over LayerNorm, which is also computationally more efficient and leads to comparable downstream performance.
@@ 5 Conclusion
We present a detailed theoretical and empirical analysis of layer normalization on hidden vectors in representation space, with a focus on the global effects of the LayerNorm operation on a vector. We first show that the LayerNorm operation can be understood in three simple steps: removing the component of a vector along the uniform vector ($\bm{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$), normalizing the remaining vector, and scaling the resultant vector by $\sqrt{d}$. We then show that LayerNorm is an irreversible process—information along the uniform vector is removed during LayerNorm and cannot be recovered using the learnable parameters available in the formulation. This is in contrast with BatchNorm, where the network has the option of learning an identity operation. We then empirically show how LayerNorm regulates the norm and the orientation of hidden vectors during model inference. Finally, we show that the “mean subtraction” operation in LayerNorm is dispensable both during inference and for training, as shown by internal representations of RMSNorm-based Llama models.
@@ Appendix A Appendix
Model Name Hidden & Uniform (Mean $\pm$ Std) Post-LN1 & Uniform (Mean $\pm$ Std) Post-LN2 & Uniform (Mean $\pm$ Std)
@@
Table 3: Average norms (L2) and standard deviations for hidden input vectors and post-layer normalization output vectors, alongside the maximum observed values for various models. ”Hidden” is an abbreviation for hidden input vectors.
Model Name Hidden (Mean $\pm$ Std) Max Hidden Post-LN1 (Mean $\pm$ Std) Max Post-LN1 Post-LN2 (Mean $\pm$ Std) Max Post-LN2
@@
Table 4: Average rotation angles in degrees between hidden and post-layer normalization vectors (LN1 and LN2) across all layers for various models, indicating the typical angular deviation introduced by normalization processes.
Model Name Hidden & Post-LN1 (Mean $\pm$ Std) Hidden & Post-LN2 (Mean $\pm$ Std)
@@
Figure 10: Distribution of angles (in degrees) between Hidden vectors (a-d) and post-normalization vectors (e-h) with the uniform vector for GPT-2 XL (Layer 43), GPT-Neo (Layer 18), Pythia 1.4 (Layer 18), and Llama-2 (Layer 24) for a layer picked at random

### v2

@@ 3 Experiments
After a detailed theoretical analysis of the effects of layer normalization on a vector, we next study how LayerNorm affects the hidden representations of LLMs in practice. While prior works have studied the effects of LayerNorm on downstream model performance and training convergence (Ba et al., 2016; Xu et al., 2019; Zhang and Sennrich, 2019), there has been a surprising gap in the literature about studying the effect of LayerNorm on the internal representations of a model at inference time. To do so, we pass one million tokens from Wikipedia articles through various models and study how LayerNorm modifies a hidden vector in the representation space.
@@
Figure 4: This figure shows the growing norm of the residual stream or hidden vectors at each layer (a-c) and how LayerNorm and RMSNorm regulate the growing norms (d-f) for GPTJ 6B, Pythia 6.9B and Llama3 8B. The dashed lines in the line plots represent one standard deviation.
Methods and Models
We study the impact of LayerNorm on the hidden representations of 7 decoder-only LLMs across two size categories. The models used are listed in Table 1. GPT2XL, GPT-Neo-1.3B, and Pythia-1.4B represent the small LLM size category, whereas GPT-J-6B, Pythia-6.9B, Llama-2-7B, and Llama-3-8B represent the medium LLM size category. We pass one million tokens from Wikipedia articles through each model and capture the hidden representations for all tokens before and after each normalization layer. This is done for each layer inside the model, which creates many terabytes of data to be analyzed. Norms were calculated using the L2 norm, and angle differences in degrees were determined through cosine similarity calculations to measure the orientation changes induced by normalization.
As outlined in Table 1, we also use two types of normalization methods. Models like GPT-2 XL, GPT-Neo 1.3B, Pythia-1.4B, GPT-J 6B, and Pythia-6.9B employ LayerNorm Ba et al. (2016), whereas Llama-2-7B and Llama-3-8B utilize RMSNorm Zhang and Sennrich (2019). The crucial difference between LayerNorm and RMSNorm is the subtraction of the mean vector (Figure 1). In this paper, we study how LayerNorm and RMSNorm operate on the hidden representations.
@@ 3.1 Norm Stabilization
As a token representation passes through the different layers inside an LLM, the hidden representations accumulate due to residual connections He et al. (2016). These hidden representations get added to the residual stream at each layer as shown in equation 17 and can cause the norm of the hidden vectors to grow. In fact, the norm of the hidden vectors at each layer grows disproportionately with large standard deviations, as can be seen in Figure 4 (a-c). The growing norm of the hidden vectors is extreme for some models more than others, as is seen for GPT-J and Pythia-6.9B. The growing norm of hidden representations for the remaining models from Table 1 and the effect of layer normalization can be seen in Figure 7 (appendix).
The significant role of LayerNorm and RMSNorm in stabilizing the growing norm of the hidden vectors at each layer can be seen in Figure 4 (d-f). The standardization step in layer normalization first modifies the norm of each vector to $\sqrt{d}$, where $d$ is the dimensionality of the representation space. The norms of the standardized vectors are further modified marginally by the scale-and-shift steps. Layer normalization affects both the mean and the spread of the norms, where the standard deviations in the norm are reduced by factors between 10-100 depending on the model (Table 3 in appendix). These results show the crucial role that LayerNorm and RMSNorm play in stabilizing the intermediate hidden vectors in practice.
@@ 3.2 Rotation
While stabilizing the growing norm of hidden vectors is a major function of layer normalization, it also ends up rotating the hidden vectors. Figure 3 shows the angle between the original hidden vector and the post-normalization vector. As can be seen in Figure 3 (and in Figure 9 for the rest of the models in the appendix), each layer is responsible for a fixed, non-trivial amount of rotation of hidden vectors in the representation space. Mean rotation angles in degrees between hidden and post-layer normalization vectors (LN1 and LN2) across all layers for various models are shown in Table 4 in the appendix. Thus, both LayerNorm and RMSNorm additionally rotate hidden vectors in representation space in addition to stabilizing the norms of the vectors.
@@
Figure 5: Distribution of angles (in degrees) between Hidden vectors (a-c) and post-normalization vectors (d-f) with the uniform vector for GPT-J, Pythia 6.9, Llama-3 for a randomly selected layer (Layer 24). The results are independent of the choice of layers.
4 LayerNorm versus RMSNorm
Apart from norm stabilization and rotation, a critical aspect of LayerNorm is its ability to orient the hidden vectors orthogonal to the uniform vector, as discussed in section 2. In section 2.2, we see that the definition of LayerNorm is innately linked with the uniform vector, possibly unintended by the original authors. Since the information along the uniform vector is being removed irreversibly during layer normalization (section 2.3), the layer normalization process implicitly assumes that the information along the uniform vector is either not important or that the model should not store information along that direction. However, two randomly chosen vectors in high-dimensional spaces are nearly orthogonal to each other with a very small spread. Since the “mean subtraction" step in LayerNorm means removing the component along the arbitrarily chosen direction of the uniform vector, the need for such a step would be justified if the hidden representations created by the model had unnaturally significant components along this vector. However, there seems to be no justification given in the literature for the mean subtraction step, including the original paper Ba et al. (2016).
RMSNorm, on the other hand, does not perform this step and simply normalizes the existing vector Zhang and Sennrich (2019). The Llama model series (Touvron et al., 2023a, b) is the most prominent family of LLMs that use RMSNorm. The simple fact that Llama models achieve state-of-the-art results across multiple measures (Dubey et al., 2024) shows that using RMSNorm instead of LayerNorm does not hurt performance. This provides a strong motivation to explore the need for the “mean subtraction" step in LayerNorm. Thus, we ask the question - do intermediate hidden representations have a non-trivial component along the uniform vector that justifies its removal?
We study this question at inference time. To answer this question, we measure the angle between the hidden vector and the uniform vector just before and after the normalization operations for LayerNorm and RMSNorm-based models. To justify the mean subtraction step, one of the following two scenarios should be true:
  • Scenario-1: In LayerNorm-based models, the intermediate hidden vectors have a large component along the uniform vector pre-normalization, which gets removed post-normalization.
  • Scenario-2: In RMSNorm-based models, which are trained without the mean subtraction step, the hidden vectors consistently have a large component along the uniform vector.
The results are shown in Figure 5 for GPT-J (6B), Pythia (6.9B), and Llama-3 (8B) and Figure 11 for the remaining models. We see that for all models in Figure 5, the angle between the hidden vectors and the uniform vectors is 90 degrees on average even before normalization, with a very small spread. This remains true post-normalization as well for all three models. This result is surprising because the hidden representations for LayerNorm-based models (GPT-J and Pythia) are themselves orthogonal to the uniform vector even before the LayerNorm operation. This is true for 4 out of the 5 LayerNorm-based models studied in this paper except for GPT2-XL (Figure 11 in the appendix). This shows that the “mean subtracting" operation of LayerNorm is redundant during inference, as there is nothing to remove.
@@ 4 LayerNorm versus RMSNorm
For RMNSNorm-based LLMs, the result for Llama-3 (8B) in Figure 5 represents scenraio -2. We see that even for Llama-3 (8B) trained using RMSNorm, the hidden representations before normalization are also on average orthogonal to the uniform vector with a very small spread. In fact, they are more orthogonal to the uniform vector than for LayerNorm-based models, with a much smaller variance. After RMSNorm, the hidden representations continue to be orthogonal to the uniform vector. This is also true for Llama-2 (7B) (Figure 11 in the appendix). This shows that even without the “mean subtraction" step during training as in RMSNorm, the hidden vectors operate orthogonal to the uniform vector, underscoring the redundance of the “mean subtraction" step.
@@
These results show that the “mean subtracting" operation in LayerNorm is redundant as the model naturally aligns its representations orthogonal to the uniform vector as expected in such high-dimensional spaces. In fact, in hindsight, evidence of having large components along the uniform vector would have been more surprising and indicated something fundamentally important about the specific vector. With this, we provide the first mechanistic evidence that the “mean subtraction" step in LayerNorm is dispensable, as we do not find hidden vectors having large components along the uniform vector. This line of investigation was made possible by our novel geometrical interpretation of LayerNorm (section 2.2), which presented the global implications of the mean subtraction step in LayerNorm - removing the component of intermediate hidden representations along the uniform vector. Due to these results, we advocate for using RMSNorm over LayerNorm, which is also computationally more efficient and leads to comparable downstream performance.
@@ 5 Conclusion
We present a detailed theoretical and empirical analysis of layer normalization on hidden vectors in representation space, with a focus on the global effects of the LayerNorm operation on a vector. We first show that the LayerNorm operation can be understood in three simple steps: removing the component of a vector along the uniform vector ($\boldsymbol{1}=[1,1,1,1,\cdots,1]^{T}\in\mathbb{R}^{d}$), normalizing the remaining vector, and scaling the resultant vector by $\sqrt{d}$. We then show that LayerNorm is an irreversible process—information along the uniform vector is removed during LayerNorm and cannot be recovered using the learnable parameters available in the formulation. This is in contrast with BatchNorm, where the network has the option of learning an identity operation. We then empirically illustrate how LayerNorm regulates the norm and the orientation of hidden vectors during model inference. Finally, we show that the “mean subtraction" operation in LayerNorm is dispensable both during inference and for training, as shown by internal representations of RMSNorm-based Llama models.
6 Limitations
This study provides valuable insights into the theoretical and empirical effects of layer normalization. Although the study empirically tests 7 models across two sizes and multiple language model families, the size of the models tested is restricted to 8 billion parameters. Since we wanted our results to have a large enough sample size, we chose to store hidden representations of 1 million tokens. Even for a small model like GPT2-XL, this requires 2TB of data, and for larger models like Llama3-8B, it requires storing approximately 4TB of data. Storing such large amounts of hidden representations pushed the capacity of the computational resources available to us. Because of this restriction, we leave testing our findings for larger models to future works and groups that have access to larger amounts of computing.
@@ Appendix A Appendix
A.1 Computations in Decoder-only LLMs
In this paper, we study the hidden representations of modern decoder-only large language models (LLMs). Let $h^{l}$ represent the intermediate hidden state vectors between each decoder layer. As depicted in Figure 6, the computations within a layer of most decoder-only LLMs proceed as follows:
[…]
$a^{l}$ $=\texttt{Att}(f^{l})$ (13)
[…]
$m^{l}$ $=W^{l}_{proj}\sigma(W^{l}_{fc}g^{l}+b^{l}_{fc})+b_{proj}$ (15)
$h^{l}$ $=h^{l-1}+a^{l}+m^{l}$ (16)
$h^{l}=h^{0}+\sum^{i=l}_{i=0}a^{i}+\sum^{i=l}_{i=0}m^{i}$ (17)
Figure 6: A high-level diagram representing the computation within one decoder block of an LLM.
A.2 Hyperparameters and Computation Resources
@@
A.3 Additional Ablations
Model Hidden & Uniform Post-LN1 & Uniform Post-LN2 & Uniform
Name (Mean $\pm$ Std) (Mean $\pm$ Std) (Mean $\pm$ Std)
@@ A.3 Additional Ablations
Model Hidden Max Post-LN1 Max Post-LN2 Max
Name (Mean $\pm$ Std) Hidden (Mean $\pm$ Std) Post-LN1 (Mean $\pm$ Std) Post-LN2
@@
Table 3: This table presents the average norms (L2) and standard deviations for hidden input vectors and post-layer normalization output vectors, alongside the maximum observed values for various models. "Hidden" is an abbreviation for hidden input vectors.
Model Hidden & Post-LN1 Hidden & Post-LN2
Name (Mean $\pm$ Std) (Mean $\pm$ Std)
@@
Table 4: This table shows the average rotation angles in degrees between hidden and post-layer normalization vectors (LN1 and LN2) across all layers for various models, indicating the typical angular deviation introduced by normalization processes.
@@
Figure 11: Distribution of angles (in degrees) between Hidden vectors (a-d) and post-normalization vectors (e-h) with the uniform vector for GPT-2 XL (Layer 43), GPT-Neo (Layer 18), Pythia 1.4 (Layer 18), and Llama-2 (Layer 24) for a specific layer chosen
(a) Pre-LN2 Pythia 6.9
(b) Pre-LN2 Llama-3
(c) Pre-LN2 GPT-2 XL
(d) Pre-LN2 GPT-Neo
(e) Pre-LN2 Pythia 1.4
(f) Pre-LN2 Llama-2
Figure 12: Distribution of angles (in degrees) between (PreLN2 and Uniform) vectors for all models for a single layer chosen (Layer 43 for GPT-2 XL, Layer 18 for GPT-Neo and Pythia 1.4, and Layer 24 for Llama-2, Pythia 6.9, and Llama-3)